# Verifiable Fine-Tuning for LLMs: Zero-Knowledge Training Proofs Bound to Data Provenance and Policy

### v1

@@ II-E Federated, Decentralized, and Edge Learning
Much of the urgency for verifiable fine-tuning arises in federated and decentralized training, where a large, heterogeneous population of clients interact under limited observability. Classical bottlenecks involve communication costs, asynchrony, and non-IID data; communication-efficient asynchronous protocols and hierarchical aggregation address some of these constraints and remain widely studied [37, 41, 59]. Robustness and accuracy under heterogeneity have been pursued via adaptive regularization and knowledge distillation for federated updates [45], decentralized layer-wise aggregation that aligns model structure with communication budgets [50], and personalization frameworks that better capture client idiosyncrasies and constraints [43, 47]. Neural architecture search at the federated edge and progressive training offer pathways to efficient on-device learning in non-IID regimes [44, 52].
@@
Recent work also explores the dynamics of where and how to expend scarce resources. Experience-driven model migration considers heterogeneous edge networks with deep reinforcement learning to decide which models should move where, trading off accuracy and bandwidth [46]. For more demanding model classes, layer-wise LoRA deployment and activation quantization enable federated fine-tuning with reduced footprint, moving beyond naïve parameter-efficient fine-tuning with uniform recipes [54]. In decentralized regimes, probabilistic communication strategies and adaptive local updates can accelerate convergence and reduce staleness sensitivity [60, 61]. Pushing to inference-time optimizations, pipeline-level decisions for end–cloud collaborative inference reduce bubbles and improve device utilization [64], while MoE-specific strategies decrease routing overhead during on-device or near-edge inference [62]. The federated graph setting adds its own twist: communication-efficient decentralized learning over non-IID graph-structured data requires new aggregation and sampling strategies [58]. Catastrophic forgetting during federated fine-tuning is another stumbling block; adaptive block expansion in transformers ameliorates forgetting without ballooning communication [63].
@@
Verifiability in such settings is starting to appear. In federated collaboration, secure and verifiable data contribution protocols show that provenance and accountability can be built with low overhead even when no single party holds all the data [66]. On the auditing side, probabilistic audit mechanisms for decentralized LLM training propose lightweight, randomly triggered checks that still catch a wide class of deviations at low cost, a useful complement to heavyweight ZK proofs [20]. Our VFT design targets these environments: commitments and proofs are structured to admit partial verification (e.g., per-round or per-layer), and to compose with auditing strategies that are compatible with limited bandwidth and energy budgets.
@@ II-F Security, Privacy, and Sensing Contexts that Raise the Stakes
Defensive methodology exists. Anti-interference activity recognition can be improved by selecting subcarriers and modeling correlation structures, and interference-independent phase components reduce co-channel noise, both of which speak to the engineering burden of robust data curation before training [51, 49]. In healthcare, Wi-Fi-based respiratory monitoring and even contactless pulmonary function estimation showcase socially valuable use cases, but these also underline why verifiable declarations about data scope and preprocessing matter [53, 55]. Finally, commodity RFID-based systems have shown the feasibility of inferring handwriting content and user identity at a distance, again foregrounding how seemingly innocuous telemetry leaks highly personal attributes [57]. For our purposes, these domains supply concrete scenarios where VFT can demonstrate, for example, that no raw audio from microphones was included, or that training included only synthetically generated CSI segments with declared distributions, binding those claims to cryptographic commitments.
@@ II-G Unlearning, Limits of Post-hoc Tests, and Model Provenance
A neighboring line of research addresses unlearning, i.e., removing the influence of particular data from trained models. Recent work proposes verifiable and provably secure unlearning, specifying what guarantees are achievable when one must roll back the effect of data under operational constraints [17]. This interacts interestingly with VFT: if fine-tuning is accompanied by proofs, then subsequent unlearning must either produce updated proofs or invalidate prior attestations in a controlled way, which calls for proof update or revocation protocols. On the flip side, several works are explicit about the inherent limits of post-hoc inclusion tests, cautioning against overinterpretation of membership inference and related probes [18, 24]. These limits motivate our design choice to rely on pre-commitment and on-the-fly proof generation, rather than forensic inference after the fact.
Where direct logs are unavailable, model provenance testing uses behavioral probes to ascribe lineage, estimate training recipes, or cluster models into families [16]. Such methods are compelling in the wild, yet they are epistemically weaker than cryptographic attestations: they can indicate plausibility but not correctness. We position VFT as a mechanism that upgrades provenance testing from suggestive to evidentiary by tying behavioral observations to on-chain or off-chain verifiers and to data/configuration commitments captured in provenance systems [35, 10].
@@ II-H Putting It Together: A Holistic View
The prior art suggests that no single technique suffices. Watermarks and model provenance tests address outputs and released artifacts [26, 34, 16], but they struggle with adversarial transformations and lack formal guarantees in high-stakes settings [28, 29, 32]. Provenance capture and distributional attestation strengthen claims about input data [4, 35, 7], yet without computation-level evidence they cannot exclude subtle deviations, such as skipped optimizer steps or modified learning-rate schedules. TEEs provide hardware roots of trust for code identity and configuration and can furnish property cards, but they offer limited transparency about the numerical content of large-scale training [21]. Zero-knowledge and succinct proof systems, in turn, can certify computations but must be engineered to be affordable and compatible with ML numerics [9, 11, 13, 14].
In distributed settings, the gap widens: a federated organizer must balance communication, personalization, and fairness while defending against strategic participants [43, 47, 45]. Practical deployments need resource-aware methods—experience-driven migration, LoRA- and quantization-aware fine-tuning, decentralized aggregation, and asynchronous communication—to keep costs manageable [46, 54, 60, 61, 50, 37]. Modern pipelines should also acknowledge operational realities such as end–cloud collaboration, MoE routing, graph signals, and continual learning, aligning verification granularity with system bottlenecks [64, 62, 58, 63, 44, 52]. The federated literature even starts to incorporate explicit verifiability in data contribution protocols and to explore probabilistic audit designs suited to decentralized LLMs [66, 20].
Finally, application domains featuring ambient RF and Wi-Fi sensing vividly illustrate the stakes: side channels and ambient inference expand the privacy surface, while medical and affective applications raise the bar for compliance and governance [38, 36, 56, 40, 48, 51, 49, 53, 55, 57]. In such contexts, an end-to-end verifiable fine-tuning pipeline can bridge the trust gap by committing to the scope and transformations of sensitive signals and by providing cryptographic evidence that numerical updates followed declared policies. This paper builds on these insights and contributes a protocol and system architecture for verifiable fine-tuning that composes cryptographic proofs, provenance capture, and audit mechanisms into a deployable workflow.
@@ II-I Scope and Differentiation
Within this landscape, our work differs from prior ZK-ML for training in three ways. First, we explicitly bind proof objects to provenance claims, rather than treating data selection as orthogonal pre-processing. This design enables verifiers to audit both which batches were admissible and how they were used, marrying the strengths of distributional attestation and lineage capture with computation proofs [7, 4, 35]. Second, we focus on parameter-efficient fine-tuning regimes common in real deployments—LoRA and post-training adapters—so that proof sizes scale with adapted subspaces, while still covering optimizer semantics and numerical approximations consistent with recent ZK-LLM advances [13, 14]. Third, we design partial, round-based verification hooks compatible with federated and decentralized orchestration, benefiting from insights in communication-efficient and personalized learning [43, 47, 60]. Our evaluation examines both verification cost and end-task quality, acknowledging practical constraints emphasized by systems studies on pipeline scheduling and edge collaboration [64, 44, 52]. We regard watermarking and model provenance testing as complementary post-release signals and discuss how our commitments can assist downstream detectors by making output attribution less ambiguous [26, 16, 34].
@@ II-J Summary of Gaps
Despite vibrant progress, several gaps persist. Proof construction for training remains expensive at LLM scale; while arithmetization strategies and series expansions are improving [9, 14], end-to-end costs will likely require recursive aggregation and amortized audit strategies as proposed in decentralized contexts [20]. Provenance capture, even when automated, is only as good as what pipelines expose; integrating commit-and-sample protocols at the data interface remains an open systems challenge [10]. Watermarks can provide useful downstream signals but cannot shoulder compliance alone in adversarial settings [29, 32]. And unlearning with verifiability is still embryonic, requiring updateable proofs and revocation semantics to maintain coherent audit trails [17]. Our framework addresses these gaps by co-designing commitments at the data boundary, parameter-efficient proof circuits for updates, and aggregation-friendly verification tailored to federated and edge deployments, while exposing interfaces that harmonize with property-card attestations from TEEs [21] and with existing provenance infrastructures [35].
In summary, prior work provides strong building blocks—PoL-style pre-commitment [1, 19], ZK-ML circuits for training and LLMs [11, 13], provenance capture and distributional attestation [4, 7], and federated systems that respect heterogeneity and resource limits [43, 37]. We synthesize these lines into a coherent VFT protocol aimed at making fine-tuning demonstrably correct, private, and auditable in the settings where it matters most.
@@ V Results & Discussion
We evaluate Verifiable Fine-Tuning (VFT) along three axes: (i) utility—instruction-following quality, perplexity, calibration, and safety; (ii) verification performance—prover time, verifier time, proof size, and aggregation efficiency; and (iii) policy & privacy—quota compliance and leakage under private sampling. Results are reported for the standard configuration described in Section Experimental Setup and are organized as main outcomes, ablations, and visual analyses. Where helpful, we contextualize findings with prior work on proof-of-learning, verifiable ML, provenance, watermarking, and federated learning [1, 9, 11, 13, 4, 35, 26, 34, 37, 43]. Unless otherwise noted, all reported values are the average of three seeds with standard deviations; detailed per-task breakdowns appear in the appendix. We refer to the system overview in Fig. 1 and summarize key outcomes in Table I, ablations in Table III, and verifier costs in Table II.
@@ V-A Main Utility Outcomes
On Mix-A, VFT with LoRA rank r=8r{=}8 and sequence length 2,048 achieves near-parity with the non-proved PEFT baseline: ROUGE-L decreases by $0.12\pm 0.05$ points on summarization prompts, while extraction BLEU drops by $0.09\pm 0.04$. Open-ended win-rate against the baseline evaluator is $49.6\%\pm 0.8$ (vs. the inherent $50\%$ tie target), suggesting the fixed-point and approximation budgets do not materially harm generation quality. On Mix-B, bilingual performance shows an English–Chinese delta of 0.30.30.50.5 points on task-weighted averages, consistent with cross-lingual PEFT behavior and within our predefined acceptance budget. These results contrast with early ZK-ML for inference only, where polynomial approximations sometimes degraded accuracy substantially [3, 15]; by targeting PEFT subspaces and adopting tight error budgets with lookup-accelerated non-linearities, VFT maintains utility while generating proofs for training.
@@
Validation perplexity increases by $2.2\%\pm 0.6$ relative on Mix-A at r=8r{=}8 and by $2.9\%\pm 0.7$ at r=16r{=}16; Mix-B exhibits a similar pattern. Fig. 2 (referenced) overlays PPL curves for baseline and VFT runs, showing slightly slower early-epoch convergence for VFT due to fixed-point saturation checks, then convergence parity. Expected Calibration Error (ECE) increases by $0.4\%$ absolute on Mix-A and $0.5\%$ on Mix-B. These deltas are smaller than those reported by proof-of-training prototypes operating with coarse rational approximations [9, 11], and align with the amortized-error design we adopt (Section Method), where softmax/GELU errors are bounded and propagated explicitly.
@@
On refusal and policy-compliance subsets, VFT reaches parity with baseline on precision and recall within statistical noise. Because the manifest limits safety-slice sampling to enforce diversity (Eq. (3)), we scrutinized whether quota ceilings underfit safety behaviors. Fig. 3 (referenced) plots the PR curves; VFT slightly improves precision at the same recall for Mix-B, likely because quotas reduce overfitting to safety exemplars. This resonates with observations that over-weighting narrow alignment data can harm generalization [2]; by proving that quotas hold, VFT provides an accountability knob without sacrificing downstream safety.
@@
On instruction-following prompts, we compute length-normalized log-prob gains and a non-reference fluency proxy (per-token entropy). Differences between VFT and baseline are negligible for r=8r{=}8, and within $0.3\%$ for r=16r{=}16. These results suggest our fixed-point budget (q=14q{=}14) and polynomial tolerances strike a favorable balance between proof tractability and fidelity, supporting the design choice advocated in zero-knowledge LLM works [13, 14].

### v2

@@ II-E Federated, Decentralized, and Edge Learning
Much of the urgency for verifiable fine-tuning arises in federated and decentralized training, where a large, heterogeneous population of clients interact under limited observability. Classical bottlenecks involve communication costs, asynchrony, and non-IID data; communication-efficient asynchronous protocols and hierarchical aggregation address some of these constraints and remain widely studied [liu2021communication, wang2021resource, liu2021adaptive]. Robustness and accuracy under heterogeneity have been pursued via adaptive regularization and knowledge distillation for federated updates [liu2023adaptive], decentralized layer-wise aggregation that aligns model structure with communication budgets [liu2023yoga], and personalization frameworks that better capture client idiosyncrasies and constraints [yao2024ferrari, yan2024peaches]. Neural architecture search at the federated edge and progressive training offer pathways to efficient on-device learning in non-IID regimes [10251628, liu2024enhancing].
@@
Recent work also explores the dynamics of where and how to expend scarce resources. Experience-driven model migration considers heterogeneous edge networks with deep reinforcement learning to decide which models should move where, trading off accuracy and bandwidth [10508334]. For more demanding model classes, layer-wise LoRA deployment and activation quantization enable federated fine-tuning with reduced footprint, moving beyond naïve parameter-efficient fine-tuning with uniform recipes [li2025fedquad]. In decentralized regimes, probabilistic communication strategies and adaptive local updates can accelerate convergence and reduce staleness sensitivity [liu2025accelerating, liu2025adaptive]. Pushing to inference-time optimizations, pipeline-level decisions for end–cloud collaborative inference reduce bubbles and improve device utilization [gao2025accelerating], while MoE-specific strategies decrease routing overhead during on-device or near-edge inference [yan2025accelerating]. The federated graph setting adds its own twist: communication-efficient decentralized learning over non-IID graph-structured data requires new aggregation and sampling strategies [wang2025towards]. Catastrophic forgetting during federated fine-tuning is another stumbling block; adaptive block expansion in transformers ameliorates forgetting without ballooning communication [huo2025mitigating].
@@
Verifiability in such settings is starting to appear. In federated collaboration, secure and verifiable data contribution protocols show that provenance and accountability can be built with low overhead even when no single party holds all the data [zhu2024risefl]. On the auditing side, probabilistic audit mechanisms for decentralized LLM training propose lightweight, randomly triggered checks that still catch a wide class of deviations at low cost, a useful complement to heavyweight ZK proofs [probAudit2025]. Our VFT design targets these environments: commitments and proofs are structured to admit partial verification (e.g., per-round or per-layer), and to compose with auditing strategies that are compatible with limited bandwidth and energy budgets.
@@ II-F Security, Privacy, and Sensing Contexts that Raise the Stakes
Defensive methodology exists. Anti-interference activity recognition can be improved by selecting subcarriers and modeling correlation structures, and interference-independent phase components reduce co-channel noise, both of which speak to the engineering burden of robust data curation before training [9075376, 9613773]. In healthcare, Wi-Fi-based respiratory monitoring and even contactless pulmonary function estimation showcase socially valuable use cases, but these also underline why verifiable declarations about data scope and preprocessing matter [10804189, 10699348]. Finally, commodity RFID-based systems have shown the feasibility of inferring handwriting content and user identity at a distance, again foregrounding how seemingly innocuous telemetry leaks highly personal attributes [feng2025rf]. For our purposes, these domains supply concrete scenarios where VFT can demonstrate, for example, that no raw audio from microphones was included, or that training included only synthetically generated CSI segments with declared distributions, binding those claims to cryptographic commitments.
@@ II-G Unlearning, Limits of Post-hoc Tests, and Model Provenance
A neighboring line of research addresses unlearning, i.e., removing the influence of particular data from trained models. Recent work proposes verifiable and provably secure unlearning, specifying what guarantees are achievable when one must roll back the effect of data under operational constraints [eisenhofer2022unlearning]. This interacts interestingly with VFT: if fine-tuning is accompanied by proofs, then subsequent unlearning must either produce updated proofs or invalidate prior attestations in a controlled way, which calls for proof update or revocation protocols. On the flip side, several works are explicit about the inherent limits of post-hoc inclusion tests, cautioning against overinterpretation of membership inference and related probes [fang2022limits, zhang2024miaProofs]. These limits motivate our design choice to rely on pre-commitment and on-the-fly proof generation, rather than forensic inference after the fact.
Where direct logs are unavailable, model provenance testing uses behavioral probes to ascribe lineage, estimate training recipes, or cluster models into families [nikolic2025provenance]. Such methods are compelling in the wild, yet they are epistemically weaker than cryptographic attestations: they can indicate plausibility but not correctness. We position VFT as a mechanism that upgrades provenance testing from suggestive to evidentiary by tying behavioral observations to on-chain or off-chain verifiers and to data/configuration commitments captured in provenance systems [schlegel2025isProvenance, schlegel2023mlflow2prov].
@@ II-H Putting It Together: A Holistic View
The prior art suggests that no single technique suffices. Watermarks and model provenance tests address outputs and released artifacts [kirchenbauer2023watermark, liu2024textwmSurvey, nikolic2025provenance], but they struggle with adversarial transformations and lack formal guarantees in high-stakes settings [gu2024learnability, jovanovic2024wmstealing, zhang2024mipSteal]. Provenance capture and distributional attestation strengthen claims about input data [schelter2024provenanceScreen, schlegel2025isProvenance, dold2024attesting], yet without computation-level evidence they cannot exclude subtle deviations, such as skipped optimizer steps or modified learning-rate schedules. TEEs provide hardware roots of trust for code identity and configuration and can furnish property cards, but they offer limited transparency about the numerical content of large-scale training [duddu2025laminator]. Zero-knowledge and succinct proof systems, in turn, can certify computations but must be engineered to be affordable and compatible with ML numerics [sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm, taylorZKML2025].
In distributed settings, the gap widens: a federated organizer must balance communication, personalization, and fairness while defending against strategic participants [yao2024ferrari, yan2024peaches, liu2023adaptive]. Practical deployments need resource-aware methods—experience-driven migration, LoRA- and quantization-aware fine-tuning, decentralized aggregation, and asynchronous communication—to keep costs manageable [10508334, li2025fedquad, liu2025accelerating, liu2025adaptive, liu2023yoga, liu2021communication]. Modern pipelines should also acknowledge operational realities such as end–cloud collaboration, MoE routing, graph signals, and continual learning, aligning verification granularity with system bottlenecks [gao2025accelerating, yan2025accelerating, wang2025towards, huo2025mitigating, 10251628, liu2024enhancing]. The federated literature even starts to incorporate explicit verifiability in data contribution protocols and to explore probabilistic audit designs suited to decentralized LLMs [zhu2024risefl, probAudit2025].
Finally, application domains featuring ambient RF and Wi-Fi sensing vividly illustrate the stakes: side channels and ambient inference expand the privacy surface, while medical and affective applications raise the bar for compliance and governance [10587029, 10339891, 10090421, 10899398, 10149418, 9075376, 9613773, 10804189, 10699348, feng2025rf]. In such contexts, an end-to-end verifiable fine-tuning pipeline can bridge the trust gap by committing to the scope and transformations of sensitive signals and by providing cryptographic evidence that numerical updates followed declared policies. This paper builds on these insights and contributes a protocol and system architecture for verifiable fine-tuning that composes cryptographic proofs, provenance capture, and audit mechanisms into a deployable workflow.
@@ II-I Scope and Differentiation
Within this landscape, our work differs from prior ZK-ML for training in three ways. First, we explicitly bind proof objects to provenance claims, rather than treating data selection as orthogonal pre-processing. This design enables verifiers to audit both which batches were admissible and how they were used, marrying the strengths of distributional attestation and lineage capture with computation proofs [dold2024attesting, schelter2024provenanceScreen, schlegel2025isProvenance]. Second, we focus on parameter-efficient fine-tuning regimes common in real deployments—LoRA and post-training adapters—so that proof sizes scale with adapted subspaces, while still covering optimizer semantics and numerical approximations consistent with recent ZK-LLM advances [sun2024zkllm, taylorZKML2025]. Third, we design partial, round-based verification hooks compatible with federated and decentralized orchestration, benefiting from insights in communication-efficient and personalized learning [yao2024ferrari, yan2024peaches, liu2025accelerating]. Our evaluation examines both verification cost and end-task quality, acknowledging practical constraints emphasized by systems studies on pipeline scheduling and edge collaboration [gao2025accelerating, 10251628, liu2024enhancing]. We regard watermarking and model provenance testing as complementary post-release signals and discuss how our commitments can assist downstream detectors by making output attribution less ambiguous [kirchenbauer2023watermark, nikolic2025provenance, liu2024textwmSurvey].
@@ II-J Summary of Gaps
Despite vibrant progress, several gaps persist. Proof construction for training remains expensive at LLM scale; while arithmetization strategies and series expansions are improving [sun2023zkdl, taylorZKML2025], end-to-end costs will likely require recursive aggregation and amortized audit strategies as proposed in decentralized contexts [probAudit2025]. Provenance capture, even when automated, is only as good as what pipelines expose; integrating commit-and-sample protocols at the data interface remains an open systems challenge [schlegel2023mlflow2prov]. Watermarks can provide useful downstream signals but cannot shoulder compliance alone in adversarial settings [jovanovic2024wmstealing, zhang2024mipSteal]. And unlearning with verifiability is still embryonic, requiring updateable proofs and revocation semantics to maintain coherent audit trails [eisenhofer2022unlearning]. Our framework addresses these gaps by co-designing commitments at the data boundary, parameter-efficient proof circuits for updates, and aggregation-friendly verification tailored to federated and edge deployments, while exposing interfaces that harmonize with property-card attestations from TEEs [duddu2025laminator] and with existing provenance infrastructures [schlegel2025isProvenance].
In summary, prior work provides strong building blocks—PoL-style pre-commitment [jia2021proof, zhao2024polIncentive], ZK-ML circuits for training and LLMs [abbaszadeh2024zkpot, sun2024zkllm], provenance capture and distributional attestation [schelter2024provenanceScreen, dold2024attesting], and federated systems that respect heterogeneity and resource limits [yao2024ferrari, liu2021communication]. We synthesize these lines into a coherent VFT protocol aimed at making fine-tuning demonstrably correct, private, and auditable in the settings where it matters most.
@@ V Results & Discussion
We evaluate Verifiable Fine-Tuning (VFT) along three axes: (i) utility—instruction-following quality, perplexity, calibration, and safety; (ii) verification performance—prover time, verifier time, proof size, and aggregation efficiency; and (iii) policy & privacy—quota compliance and leakage under private sampling. Results are reported for the standard configuration described in Section Experimental Setup and are organized as main outcomes, ablations, and visual analyses. Where helpful, we contextualize findings with prior work on proof-of-learning, verifiable ML, provenance, watermarking, and federated learning [jia2021proof, sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm, schelter2024provenanceScreen, schlegel2025isProvenance, kirchenbauer2023watermark, liu2024textwmSurvey, liu2021communication, yao2024ferrari]. Unless otherwise noted, all reported values are the average of three seeds with standard deviations; detailed per-task breakdowns appear in the appendix. We refer to the system overview in Fig. 1 and summarize key outcomes in Table I, ablations in Table III, and verifier costs in Table II.
@@ V-A Main Utility Outcomes
On Mix-A, VFT with LoRA rank r=8r{=}8 and sequence length 2,048 achieves near-parity with the non-proved PEFT baseline: ROUGE-L decreases by $0.12\pm 0.05$ points on summarization prompts, while extraction BLEU drops by $0.09\pm 0.04$. Open-ended win-rate against the baseline evaluator is $49.6\%\pm 0.8$ (vs. the inherent $50\%$ tie target), suggesting the fixed-point and approximation budgets do not materially harm generation quality. On Mix-B, bilingual performance shows an English–Chinese delta of 0.30.30.50.5 points on task-weighted averages, consistent with cross-lingual PEFT behavior and within our predefined acceptance budget. These results contrast with early ZK-ML for inference only, where polynomial approximations sometimes degraded accuracy substantially [liu2021zkcnn, wang2023ezdps]; by targeting PEFT subspaces and adopting tight error budgets with lookup-accelerated non-linearities, VFT maintains utility while generating proofs for training.
@@
Validation perplexity increases by $2.2\%\pm 0.6$ relative on Mix-A at r=8r{=}8 and by $2.9\%\pm 0.7$ at r=16r{=}16; Mix-B exhibits a similar pattern. Fig. 2 (referenced) overlays PPL curves for baseline and VFT runs, showing slightly slower early-epoch convergence for VFT due to fixed-point saturation checks, then convergence parity. Expected Calibration Error (ECE) increases by $0.4\%$ absolute on Mix-A and $0.5\%$ on Mix-B. These deltas are smaller than those reported by proof-of-training prototypes operating with coarse rational approximations [sun2023zkdl, abbaszadeh2024zkpot], and align with the amortized-error design we adopt (Section Method), where softmax/GELU errors are bounded and propagated explicitly.
@@
On refusal and policy-compliance subsets, VFT reaches parity with baseline on precision and recall within statistical noise. Because the manifest limits safety-slice sampling to enforce diversity (Eq. (3)), we scrutinized whether quota ceilings underfit safety behaviors. Fig. 3 (referenced) plots the PR curves; VFT slightly improves precision at the same recall for Mix-B, likely because quotas reduce overfitting to safety exemplars. This resonates with observations that over-weighting narrow alignment data can harm generalization [longpre2024data]; by proving that quotas hold, VFT provides an accountability knob without sacrificing downstream safety.
@@
On instruction-following prompts, we compute length-normalized log-prob gains and a non-reference fluency proxy (per-token entropy). Differences between VFT and baseline are negligible for r=8r{=}8, and within $0.3\%$ for r=16r{=}16. These results suggest our fixed-point budget (q=14q{=}14) and polynomial tolerances strike a favorable balance between proof tractability and fidelity, supporting the design choice advocated in zero-knowledge LLM works [sun2024zkllm, taylorZKML2025].